A patient (72kg) is treated with a diuretic drug. The drug is available as tablets with oral bioavailability of 67%. if the Vd is 1.8L/kg and the desired concentration is 1.4mg/L, which of the following is the most appropriate oral loading dose for this diuretic?

270 mg

408 mg

18 mg

.3 mg

620 mg
Answer #1

A patient (72kg) is treated with a diuretic drug. The drug is available as tablets with oral bioavailability of 67%. if the Vd is 1.8L/kg and the desired concentration is 1.4mg/L.

Formula to calculate dose is

Amount of Drug = Desired Conc \times Volume

Then, Total volume is 72kg x 1.8L/kg = 129.6L

Then the Amount of drug needed is

1.4mg/L x 129.6L = 181.4mg

Now, Oral bioavailability of this drug is 67%

So, 181,4 - 67 x 100 = 270.7 Mg (approx 270 mg)

Answer - 270mg